There are 8.340 x 1025 I2 molecules and 4.76 x 1030 F2 molecules in a container. These are made to react and form IF7 molecules. Which is the limiting reactant (LR) and which is the excess reactant (ER)? Use factor label method for your calculations.

Select one:

a. LR: I; ER: F

b. LR: F; ER: I

c. LR: F2; ER: I2

d. LR: I2; ER: F2
Answer #1

Reaction is

I2 + 7 F2 ...............> 2 IF7

In this reaction, 1 mole I2 react with 7 mole F2 to form 2 mole IF7.

we have

8.340 x 1025 I2 molecules and 4.76 x 1030 F2 molecules

mole of I2 molecules = 8.340 x 1025 / 6.023 * 10^23 = 138.47 mole.

and

mole of F2 = 4.76 x 1030 / 6.023 * 10^23 = 7.903 * 10^6 mole.

138.47 mole I2 should react with 138.47 * 7 = 969.29 mole F2 but there are 7.903 * 10^6 mole F2.

so

I2 is the limiting reactant (LR)

and

F2 remain in excess (ER).

option d. LR: I2; ER: F2 is the answer.